A Brayton cycle operates with air with minimum and maximum temperatures of 27 ºC and 727 ºC. It is designed so that the maximum cycle pressure is 2000 kPa and the minimum cycle temperature is 100 kPa. The isentropic efficiencies of the turbine and compressor are both 90 percent. Using constant room-temperature specific heats, (a) determine the net work produced per unit mass of air each time this cycle is executed, in kJ/kg. (b) Determine the cycle's thermal efficiency. Remember that the cycle is not an ideal cycle.
Answer #1

T1=276 300K T3 = 727°C = low K P2 - Pes. P3P: - 2000 - 20 Pi P Pus P 10 M = 0.9 = Tas-Ti acil T-TI Mp=0.g - Torty - i) To-TusPlease dear thumpsup for the answer. Thank you.